OpenStudy (anonymous):

The perimeter of a rectangle is 350ft and area 7500ft. What are the dimensions of the rectangle?

OpenStudy (tkhunny):

Area = Length * Width Perimeter = 2*Length + 2*Width

OpenStudy (anonymous):

 This is a system of two equations, isn't it? xy = 7500 2x + 2y = 350 Can you go from there?

OpenStudy (anonymous):

so far i have 350/2 = L + W then L= 175 -W

OpenStudy (anonymous):

Yep, that is right. Now, you can substitute the value you found for L (175-W) into the other equation (L*W = 7500) and solve for W